# Neural Value Iteration

### v1

@@ 1 Introduction
The partially observable Markov decision process (POMDP) framework models decision-making problems where the true state is hidden and transitions are stochastic. In a POMDP, the agent maintains a belief (a probability distribution over states) and seeks an optimal policy mapping beliefs to actions to maximize expected return. However, solving POMDPs optimally is computationally intractable: PSPACE-complete in the finite-horizon case (papadimitriou1987complexity) and undecidable for infinite horizons (madani1999undecidability).
It has been proved that the POMDP’s optimal value function $V^{*}_{t}$ at each time step $t$ is piecewise linear and convex (PWLC) (sondik1971optimal). As a result, $V^{*}_{t}$ can be efficiently represented by a finite set of $|S|$-dimensional hyperplanes, known as $\alpha$-vectors, where $S$ denotes the state space. By leveraging this PWLC structure, value iteration methods for POMDPs (e.g., dynamic programming updates) can be performed by iteratively maintaining and updating a finite set of $\alpha$-vectors (Pineau-ijcai03; Smith-uai04; Smith_HSVI2; sarsop; bai2010monte). However, since each $\alpha$-vector spans the entire state space, these methods are difficult to scale due to memory and computation costs. Even efficient solvers like SARSOP (sarsop) struggle with domains containing millions of states, where optimizing large sets of $\alpha$-vectors becomes intractable. To improve scalability in continuous-state POMDPs, MCVI (bai2010monte) implicitly represents $\alpha$-vectors using nodes in a policy graph (i.e., a finite-state controller) and performs approximate Bellman backups via Monte Carlo (MC) simulations. Despite its scalability advantages, MCVI suffers from performance degradation in large problems with long horizons due to the high cost of MC simulations, limiting its practicality unless techniques like macro-actions (macro_action_mcvi) are used to reduce planning depth and computational overhead.
@@ 2 Related Work
POMDP planning techniques can be broadly categorized into offline and online methods. Offline solvers aim to compute a complete solution prior to execution, typically in the form of a compact policy structure that can be readily executed at runtime. In contrast, online planners interleave planning and execution by computing the best action for the current belief at each decision step, often using lookahead search or sampling techniques. In this section, we mainly review the offline POMDP approaches that are most relevant to our contributions. For comprehensive surveys and broader discussions, we refer the reader to existing literature (lauri2022partially; kurniawati2022partially).
@@
Most offline POMDP solvers (Smith-uai04; Smith_HSVI2; sarsop) rely on iteratively updating the value function via Bellman backups on a finite set of $\alpha$-vectors, which represent the POMDP value function. All these methods and variants (bai2010monte; macro_action_mcvi) can be viewed as following a bottom-up approach: to solve a POMDP of horizon $h$, they recursively construct the optimal value function backward in time, from $t=h$ to $t=1$. Modern methods such as HSVI and SARSOP show that it is efficient to perform Bellman backups over a relatively small belief space by only sampling representative beliefs via reachability heuristics, which largely speeds up value iteration convergence. However, for very large POMDPs, even these value iteration methods are considered intractable. An exception among recent offline methods is POMCGS (you2025), which does not follow a value iteration and instead uses a top-down strategy inspired by online MCTS-based planners (pomcp; despot; pomcpow; adaops; UCT). It treats the POMDP as a belief-MDP and performs tree search directly in belief space, merging similar nodes by comparing belief differences across depths to compact the policy tree into a graph and reduce redundant computation. Its Monte-Carlo search framework also supports techniques such as action progressive widening, enabling efficient planning in continuous domains.
On the other hand, approximation techniques have been extensively studied to scale POMDPs to large or continuous domains. A common approach is to approximate beliefs or value functions using parametric forms, particularly Gaussian mixtures (thrun1999monte; spaan2005perseus; porta2006point; prentice2010belief). However, such representations often suffer from oversmoothing and struggle with discontinuities (e.g., walls or obstacles), requiring many Gaussian components to achieve pratical approximation, which prevents such methods from being applied on more complex domains.
@@
Similar ideas have emerged in the reinforcement learning (RL) community. In POMDP settings, RL-based methods (drqn; ppo; karkus2017qmdp) often use recurrent neural networks (RNNs) to encode a hidden state as a belief surrogate. The value function is then approximated by a neural network over this hidden representation. Recent methods like BetaZero (moss2023betazero) learn value functions over simplified belief statistics (e.g., mean and variance), using them as heuristics in MCTS planning over belief-MDPs. However, none of these approaches explicitly leverage the piecewise-linear and convex (PWLC) structure of the POMDP value function, limiting their alignment with value iteration theory and associated guarantees.
@@ 3.2 Alpha Vectors and Finite-State Controller
It is well known that a value function represented with a finite set of $\alpha$-vectors can be equivalently encoded as a finite-state controller (FSC), or policy graph, where each controller node corresponds to an $\alpha$-vector and is labeled with an action and a transition rule based on observations (hansen1997improved). This transformation enables compact policy representations and efficient execution via table lookups, especially when the size of the FSC is small relative to the belief space.
@@ 3.3 Bellman Backups in POMDPs
In this section, we describe the core Bellman backup operation that underpins our contribution. For a broader overview of point-based methods, we refer the reader to shani2013survey. A POMDP’s optimal value function can be computed through value iterations, and each iteration is commonly referred to as a Bellman Backup. This backup process updates a new value function $V^{\prime}$ from the current value function $V$ as:
@@ 6 $\alpha^{\prime}\leftarrow\arg\max_{a\in A}\left(\alpha_{a}\cdot b\right)$;
Nonetheless, when the state space is large or continuous, even this backup process becomes intractable due to the need to sum over all possible states. To address this challenge, Monte Carlo Value Iteration (MCVI) bai2010monte was proposed. MCVI avoids explicit loops over the entire state space by sampling states and implicitly representing $\alpha$-vectors using policy graph nodes. It approximates the Bellman backup using Monte Carlo simulations, a procedure referred to as MC-Backup. However, the lack of explicit $\alpha$-vector representations in MC-Backup prevents caching and reusing computation across belief points. As a result, MCVI often requires a large amount of simulation to achieve competitive performance, which limits its solution quality in large-scale problems under constrained computational budgets.
@@ 4.1 Represent $V$ with a Finite Set of Neural Networks
As discussed in the related work section, one natural approach to handling large-scale problems is to directly approximate the value function as a parametric mapping from the belief space $\mathbb{B}$ to a scalar value in $\mathbb{R}$. However, this approximation first requires the belief inputs to be some sort of simplified structure such as means and variances (moss2023betazero), or RNN’s hidden states (drqn; ppo), which may lose some information compared with the original belief. Additionally, updating the approximation for new beliefs may change its output for previously seen ones, introducing instability and reducing accuracy, especially when generalizing to unseen beliefs.
@@ 5 Experiment
We implement our main contribution, NVI, in Julia using the POMDP.jl framework (pomdpjl). Planning experiments are conducted on a machine with an Intel i7 3.6GHz CPU, 64GB RAM, and an RTX 4060 GPU, while RL experiments are performed on an NVIDIA A100-SXM4-40GB GPU.
@@ 5.1 Benchmark Domains
The Light Dark problem (pomcpow) involves an agent moving in a 1D world. The agent does not know its initial location and only receives accurate observations in a specific light area. The agent’s goal is to localize itself and navigate to a target region.
@@
The Lidar Roomba domain (roomba) features a cleaning robot that must localize itself in a 2D environment and reach a goal position. The robot has no knowledge of its initial location and is equipped with a noisy lidar sensor that measures forward distance.
@@
The RockSample (RS) problem (Smith-uai04) is defined by a tuple $\langle n,k\rangle$, where a robot moves in an $n\times n$ grid world containing $k$ rocks. Initially, the robot knows its own location but not the rocks’ status (good or bad). The robot can measure rock quality with a noisy sensor whose accuracy depends on the distance to the rock. The goal is to collect as many good rocks as possible before reaching the target location.
@@
Figure 2: NBB’s upper and lower bounds with backup numbers on RS (20,20), Light Dark, and Lidar Roomba. Baseline online planning methods (despot; pomcpow; cai2021hyp; adaops) are reported from BetaZero (moss2023betazero).
@@ 5.2 Offline POMDP Methods
Table 1: Performance comparison of different offline POMDP algorithms. – The symbol $\dagger$" indicates out-of-memory errors during computation, and “—" denotes algorithm inapplicability to the domain. All methods’ performance is evaluated over 5 different runs, with sampling-based approaches’ policies evaluated using $10^{5}$ simulations per run. In RockSample domains, each run uses a distinct configuration with randomized rock positions and statuses, while the robot always starts at position $(1,1)$ to maximize the distance to the exit area.
@@
Last but not least, despite an extensive parameter search for Recurrent PPO (as detailed in the appendix), it failed to produce competitive policies compared to offline planning methods in most domains. Notably, while some individual episodes during training showed promising behavior, the learning process frequently converged to a local optimum corresponding to a myopic solution (e.g., in RockSample, the agent moves directly to the exit area). Notably, tao2025pobax report higher performance in the RockSample problem, which may be attributed to differences in parameter settings or problem configurations.
@@ 6 Discussion
The idea of representing $\alpha$-vectors with neural networks opens several directions for future research. Techniques such as state sampling (pmlr-v28-brechtel13), belief selection heuristics, macro actions (macro_action_mcvi), and observation grouping can be incorporated into Neural Value Iteration to further improve scalability in complex and continuous domains.
@@ 7 Conclusion
In our work, we adopt a standard belief collection technique similar to HSVI Smith_HSVI2 during value iteration. The detailed procedure is provided in Algorithm˜4. We maintain a belief tree $\mathcal{T}$ and perform forward search to collect beliefs along sampled trajectories. At each step, actions are selected according to Q-values, and observations are chosen based on maximum excess uncertainty. Upper bounds are initialized using a QMDP heuristic, while lower bounds are updated through value backups.
@@ 17 $\mathcal{B}\leftarrow\mathcal{B}\cup\{b^{a,o}\}$ ;
For the RL experiments, we use the pobax package (tao2025pobax), which is entirely written in JAX (jax2018github) [^1] Commit 4e53aac is the latest version of the code we used. RL experiments require significant amounts of fine-tuning hyperparameters, which is why JAX is currently the state-of-the-art choice, since it can run multiple runs in parallel on the GPU.
@@
The network architecture of the actor critic model is the same as in tao2025pobax, which is shown in Fig. 4.
@@
Figure 4: Actor critic architecture similar as tao2025pobax
We performed hyperparameter tuning for each environment using the optuna package (optuna_2019) with 50 trials and the following ranges:

### v2

@@ 1 Introduction
The partially observable Markov decision process (POMDP) framework models decision-making problems where the true state is hidden and transitions are stochastic. In a POMDP, the agent maintains a belief (a probability distribution over states) and seeks an optimal policy mapping beliefs to actions to maximize expected return. However, solving POMDPs optimally is computationally intractable: PSPACE-complete in the finite-horizon case (Papadimitriou and Tsitsiklis, 1987) and undecidable for infinite horizons (Madani et al., 1999).
It has been proved that the POMDP’s optimal value function $V^{*}_{t}$ at each time step $t$ is piecewise linear and convex (PWLC) (Sondik, 1971). As a result, $V^{*}_{t}$ can be efficiently represented by a finite set of $|S|$-dimensional hyperplanes, known as $\alpha$-vectors, where $S$ denotes the state space. By leveraging this PWLC structure, value iteration methods for POMDPs (e.g., dynamic programming updates) can be performed by iteratively maintaining and updating a finite set of $\alpha$-vectors (Pineau et al., 2003; Smith and Simmons, 2004, 2005; Kurniawati et al., 2008; Bai et al., 2010). However, since each $\alpha$-vector spans the entire state space, these methods are difficult to scale due to memory and computation costs. Even efficient solvers like SARSOP (Kurniawati et al., 2008) struggle with domains containing millions of states, where optimizing large sets of $\alpha$-vectors becomes intractable. To improve scalability in continuous-state POMDPs, MCVI (Bai et al., 2010) implicitly represents $\alpha$-vectors using nodes in a policy graph (i.e., a finite-state controller) and performs approximate Bellman backups via Monte Carlo (MC) simulations. Despite its scalability advantages, MCVI suffers from performance degradation in large problems with long horizons due to the high cost of MC simulations, limiting its practicality unless techniques like macro-actions (Lim et al., 2011) are used to reduce planning depth and computational overhead.
@@ 2 Related Work
POMDP planning techniques can be broadly categorized into offline and online methods. Offline solvers aim to compute a complete solution prior to execution, typically in the form of a compact policy structure that can be readily executed at runtime. In contrast, online planners interleave planning and execution by computing the best action for the current belief at each decision step, often using lookahead search or sampling techniques. In this section, we mainly review the offline POMDP approaches that are most relevant to our contributions. For comprehensive surveys and broader discussions, we refer the reader to existing literature (Lauri et al., 2022; Kurniawati, 2022).
@@
Most offline POMDP solvers (Smith and Simmons, 2004, 2005; Kurniawati et al., 2008) rely on iteratively updating the value function via Bellman backups on a finite set of $\alpha$-vectors, which represent the POMDP value function. All these methods and variants (Bai et al., 2010; Lim et al., 2011) can be viewed as following a bottom-up approach: to solve a POMDP of horizon $h$, they recursively construct the optimal value function backward in time, from $t=h$ to $t=1$. Modern methods such as HSVI and SARSOP show that it is efficient to perform Bellman backups over a relatively small belief space by only sampling representative beliefs via reachability heuristics, which largely speeds up value iteration convergence. However, for very large POMDPs, even these value iteration methods are considered intractable. An exception among recent offline methods is POMCGS (You et al., 2025), which does not follow a value iteration and instead uses a top-down strategy inspired by online MCTS-based planners (Silver and Veness, 2010; Somani et al., 2013; Sunberg and Kochenderfer, 2017; Wu et al., 2021; Kocsis and Szepesvári, 2006). It treats the POMDP as a belief-MDP and performs tree search directly in belief space, merging similar nodes by comparing belief differences across depths to compact the policy tree into a graph and reduce redundant computation. Its Monte-Carlo search framework also supports techniques such as action progressive widening, enabling efficient planning in continuous domains.
@@
On the other hand, approximation techniques have been extensively studied to scale POMDPs to large or continuous domains. A common approach is to approximate beliefs or value functions using parametric forms, particularly Gaussian mixtures (Thrun, 1999; Spaan and Vlassis, 2005; Porta et al., 2006; Prentice and Roy, 2010). However, such representations often suffer from oversmoothing and struggle with discontinuities (e.g., walls or obstacles), requiring many Gaussian components to achieve pratical approximation, which prevents such methods from being applied on more complex domains. Similar ideas have emerged in the reinforcement learning (RL) community. In POMDP settings, RL-based methods (Hausknecht and Stone, 2017; Schulman et al., 2017; Karkus et al., 2017) often use recurrent neural networks (RNNs) to encode a hidden state as a belief surrogate. The value function is then approximated by a neural network over this hidden representation. Recent methods like BetaZero (Moss et al., 2023) learn value functions over simplified belief statistics (e.g., mean and variance), using them as heuristics in MCTS planning over belief-MDPs. However, none of these approaches explicitly leverage the piecewise-linear and convex (PWLC) structure of the POMDP value function, limiting their alignment with value iteration theory and associated guarantees.
@@ 3.2 Alpha Vectors and Finite-State Controller
It is well known that a value function represented with a finite set of $\alpha$-vectors can be equivalently encoded as a finite-state controller (FSC), or policy graph, where each controller node corresponds to an $\alpha$-vector and is labeled with an action and a transition rule based on observations (Hansen, 1997). This transformation enables compact policy representations and efficient execution via table lookups, especially when the size of the FSC is small relative to the belief space.
@@ 3.3 Bellman Backups in POMDPs
In this section, we describe the core Bellman backup operation that underpins our contribution. For a broader overview of point-based methods, we refer the reader to Shani et al. (2013). A POMDP’s optimal value function can be computed through value iterations, and each iteration is commonly referred to as a Bellman Backup. This backup process updates a new value function $V^{\prime}$ from the current value function $V$ as:
@@ 6 $\alpha^{\prime}\leftarrow\arg\max_{a\in A}\left(\alpha_{a}\cdot b\right)$;
Nonetheless, when the state space is large or continuous, even this backup process becomes intractable due to the need to sum over all possible states. To address this challenge, Monte Carlo Value Iteration (MCVI) Bai et al. (2010) was proposed. MCVI avoids explicit loops over the entire state space by sampling states and implicitly representing $\alpha$-vectors using policy graph nodes. It approximates the Bellman backup using Monte Carlo simulations, a procedure referred to as MC-Backup. However, the lack of explicit $\alpha$-vector representations in MC-Backup prevents caching and reusing computation across belief points. As a result, MCVI often requires a large amount of simulation to achieve competitive performance, which limits its solution quality in large-scale problems under constrained computational budgets.
@@ 4.1 Represent $V$ with a Finite Set of Neural Networks
As discussed in the related work section, one natural approach to handling large-scale problems is to directly approximate the value function as a parametric mapping from the belief space $\mathbb{B}$ to a scalar value in $\mathbb{R}$. However, this approximation first requires the belief inputs to be some sort of simplified structure such as means and variances (Moss et al., 2023), or RNN’s hidden states (Hausknecht and Stone, 2017; Schulman et al., 2017), which may lose some information compared with the original belief. Additionally, updating the approximation for new beliefs may change its output for previously seen ones, introducing instability and reducing accuracy, especially when generalizing to unseen beliefs.
@@ 4.2 Value Iteration with Neural Bellman Backups
4.3 Benefits of $\alpha$-vector Recasting
We analyze the detailed benefits of recasting $\alpha$-vectors with neural networks.
A significant advantage is the exponentially reduced number of simulator calls compared to MCVI. Specifically, in each MCVI backup, since the algorithm does not maintain an explicit $\alpha$-vector representation, the value $\alpha(s^{\prime})$ (or $V_{n,s^{\prime}}$ in Algorithm˜3) must be estimated via Monte-Carlo simulations using the current policy starting from node $n$ and state $s^{\prime}$. To obtain a reliable estimate, a large number of rollouts ($nb_{\text{rollout}}$) and considerable simulation depth ($d$) are typically required. Consequently, each MCVI backup requires looping over actions ($a\in A$), belief particles ($s\in b$), and nodes ($n\in G$), demanding approximately
$|A|\times nb_{\text{particle}}\times nb_{\text{node}}\times nb_{\text{rollout}}\times d$
simulator calls to access all $\alpha(s^{\prime})$. Even with a fast simulator, this cost remains high and grows with the size of the node set. More critically, for larger or highly stochastic problems, the required number of belief particles $nb_{\text{particle}}$, simulation depth $d$, and rollouts $nb_{\text{rollout}}$ must increase further to ensure accurate Monte‑Carlo evaluation; otherwise, performance degradation may occur as observed in Lim et al. (2011).
In contrast, like other offline solvers that maintain explicit $\alpha$-vector representations, NVI can directly access the value $\alpha(s^{\prime})$ and, more importantly, predict $\alpha(s^{\prime})$ in batches with negligible time. This capability reduces the required simulator steps by orders of magnitude compared to MCVI, yielding substantial computational savings even when the simulator is fast. One might argue that the burden is merely shifted to the GenData process and neural network training. However, as explained in the previous section, only the first NVI backup (when $G$ is empty) requires full rollouts, and those rollouts loop only over $nb_{\text{sample}}$ states—not over actions and each belief particle as in MCVI. Moreover, each NVI backup trains only a small neural network (e.g., an MLP) using $nb_{\text{sample}}\ll|S|$ samples, making the training load very light and easily accelerated on modern GPUs.
A further benefit emerges when comparing NVI to classic offline solvers such as SARSOP or HSVI, which store $\alpha$-vectors explicitly. For large problems like RockSample($20,20$) with a state space |S|>400|S|>400M, even minimal storage per vector exceeds 14GB, rendering such approaches memory‑infeasible. By contrast, a 3‑layer MLP (with 1024 neurons per layer) representing $\alpha_{\text{NN}}$ requires <20<20MB.
These observations precisely identify the scalability barriers that have limited prior offline solvers. Sampling‑based methods like MCVI can handle small‑ to medium‑scale problems, but their reliance on costly Monte‑Carlo simulations restricts their practicality in large, long‑horizon domains Lim et al. (2011); obtaining a near‑optimal policy in such settings may demand unacceptable computation time. Solvers that store $\alpha$-vectors explicitly, such as SARSOP and HSVI, can deliver $\epsilon$-optimal solutions for small and certain medium‑size problems. However, without exploiting state‑gradient information, their memory requirements for storing $\alpha$-vectors explicitly become prohibitive for large state spaces.
NVI addresses these limitations by leveraging state gradient information through the representation of $\alpha$-vectors as $\alpha_{\text{NN}}$, making $\alpha$-vectors tractable in huge state spaces with acceptable fitting error. We argue that this learning for planning approach, where neural representations preserve the state-gradient structure of $\alpha$-vectors, is a crucial step toward scalable offline POMDP planning.
@@ 5 Experiment
We implement our main contribution, NVI, in Julia using the POMDP.jl framework (Egorov et al., 2017). Planning experiments are conducted on a machine with an Intel i7 3.6GHz CPU, 64GB RAM, and an RTX 4060 GPU, while RL experiments are performed on an NVIDIA A100-SXM4-40GB GPU.
@@ 5.1 Benchmark Domains
The Light Dark problem (Sunberg and Kochenderfer, 2017) involves an agent moving in a 1D world. The agent does not know its initial location and only receives accurate observations in a specific light area. The agent’s goal is to localize itself and navigate to a target region.
@@
The Lidar Roomba domain (Menda et al., 2023) features a cleaning robot that must localize itself in a 2D environment and reach a goal position. The robot has no knowledge of its initial location and is equipped with a noisy lidar sensor that measures forward distance.
@@
The RockSample (RS) problem (Smith and Simmons, 2004) is defined by a tuple $\langle n,k\rangle$, where a robot moves in an $n\times n$ grid world containing $k$ rocks. Initially, the robot knows its own location but not the rocks’ status (good or bad). The robot can measure rock quality with a noisy sensor whose accuracy depends on the distance to the rock. The goal is to collect as many good rocks as possible before reaching the target location.
@@
Figure 2: NBB’s upper and lower bounds with backup numbers on RS (20,20), Light Dark, and Lidar Roomba. Baseline online planning methods (Somani et al., 2013; Sunberg and Kochenderfer, 2017; Cai et al., 2021; Wu et al., 2021) are reported from BetaZero (Moss et al., 2023).
@@ 5.2 Offline POMDP Methods
Table 1: Performance comparison of different offline POMDP algorithms. – The symbol $\dagger$" indicates out-of-memory errors during computation, and “" denotes algorithm inapplicability to the domain. All methods’ performance is evaluated over 5 different runs, with sampling-based approaches’ policies evaluated using $10^{5}$ simulations per run. In RockSample domains, each run uses a distinct configuration with randomized rock positions and statuses, while the robot always starts at position $(1,1)$ to maximize the distance to the exit area.
@@
Last but not least, despite an extensive parameter search for Recurrent PPO (as detailed in the appendix), it failed to produce competitive policies compared to offline planning methods in most domains. Notably, while some individual episodes during training showed promising behavior, the learning process frequently converged to a local optimum corresponding to a myopic solution (e.g., in RockSample, the agent moves directly to the exit area). Notably, Tao et al. (2025) report higher performance in the RockSample problem, which may be attributed to differences in parameter settings or problem configurations.
@@ 6 Discussion
The idea of representing $\alpha$-vectors with neural networks opens several directions for future research. Techniques such as state sampling (Brechtel et al., 2013), belief selection heuristics, macro actions (Lim et al., 2011), and observation grouping can be incorporated into Neural Value Iteration to further improve scalability in complex and continuous domains.
@@ 7 Conclusion
In our work, we adopt a standard belief collection technique similar to HSVI Smith and Simmons (2005) during value iteration. The detailed procedure is provided in Algorithm˜4. We maintain a belief tree $\mathcal{T}$ and perform forward search to collect beliefs along sampled trajectories. At each step, actions are selected according to Q-values, and observations are chosen based on maximum excess uncertainty. Upper bounds are initialized using a QMDP heuristic, while lower bounds are updated through value backups.
@@ 17 $\mathcal{B}\leftarrow\mathcal{B}\cup\{b^{a,o}\}$ ;
For the RL experiments, we use the pobax package (Tao et al., 2025), which is entirely written in JAX (Bradbury et al., 2018) [^1] Commit 4e53aac is the latest version of the code we used. RL experiments require significant amounts of fine-tuning hyperparameters, which is why JAX is currently the state-of-the-art choice, since it can run multiple runs in parallel on the GPU.
@@
The network architecture of the actor critic model is the same as in Tao et al. (2025), which is shown in Fig. 4.
@@
Figure 4: Actor critic architecture similar as Tao et al.
We performed hyperparameter tuning for each environment using the optuna package (Akiba et al., 2019) with 50 trials and the following ranges: